Displaying reports 561-580 of 1091.Go to page Start 25 26 27 28 29 30 31 32 33 End
Reports until 12:39, Tuesday 16 July 2024
GQuEST Output Filter Cavities (GQuEST)
Erin McGee - posted 12:39, Tuesday 16 July 2024 (11747)
Filter Cavity

[Erin, Torrey]

Ran a fiber with 1500nm light from the larger table to the other and changed the alignment of the BS leading into the mirror to maximize the power going to the other table (~40mW). We also added a fiber beam splitter to the filter cavity setup, and ran the fiber into the setup.

We performed a rough alignment of the mirrors, so that the light passes through both of the cavities.

Also added the three flip mirrors to the setup, aligned them, and connected buttons and power to them.

Equipment General
Torrey Cullen - posted 11:20, Tuesday 16 July 2024 - last comment - 14:00, Tuesday 16 July 2024(11746)
QuantaTech AOMs

AOMs are being stored in B102, beige cabinet on the east side, top right shelf.

Most update to date location of the AOMs can be found at  https://wiki.mccullerlab.com/DCC/MX-S24004.

Images attached to this report
Comments related to this report
Lee McCuller - 14:00, Tuesday 16 July 2024 (11748)

Please add labels with unit numbers as you pull them out.

Atoms General
Briana Chen - posted 08:01, Tuesday 16 July 2024 - last comment - 19:14, Tuesday 16 July 2024(11743)
More absorption dips/temperature control setup

[Briana, Ian]

Set up TC300 (vapor cell temperature controller), which is next to the laser controllers. Even though the manual for the TC300 vapor cell temperature controller states that using a 4-wire setting for better accuracy, it does not apply to our thermistor. We should be on the 2-wire setting. If the control isn't applying, make sure the channel buttons are illuminated in green as otherwise they're disabled. Currently having an issue with the discretization of analog temperature output from TC300 as there is only a sharp dip in voltage (corresponding to temperature) when the temperature read reaches an integer (for example, from 23.0 to 22.01 Celsius, the voltage reading is roughly constant and once 22.0 Celsius is reached, the voltage reading drops significantly). See attached picture (temp_ctrl_discrete). I wonder if the 20 Ohm impedance from the analog output needs to be matched by the impedance BNC cable connecting to it. I will also try expanding the temperature range as the voltage drop is equal to the 5V spanned by the analog output divided by the set temperature range (0-50 Celsius), so maybe that has an effect. This range was chosen because the heater assembly should not exceed 50 deg Celsius. Either way, will need to fix this before tuning PID controllers of the TC300. 

We moved the bottom Moku Pro (not connected to anything) on the rack near the main computer in place of where the Moku Go was. Moku Go has been returned to original cabinet. Both channels of the vapor cell temperature controller (labelled Aliyah Boston and Angel Reese) have been connected to the Moku input. Photodetector connected to Moku input is labelled Caitlin Clark. Moku output (connected to laser temperature controller input) is labelled Joey Votto. 

We repeated the measurements last post with the current controller set. We set the laser current controller at 125.08 mA, laser temperature controller scanning from 6.871 - 7.160 kOhms (32.8 - 33.78 Celsius), and the temperature setting on the vapor cell to 45 degrees Celsius on both controllers. Current limit is 0.7 A. We see two dips, one small and one large. The large one likely corresponds to the 780 nm transition, but maybe there is some hyperfine structure causing the smaller dip. We can determine the temperature at which these peaks occur and then, using laser specs, the wavelength, which we can then use to deduce the transition (need to call Thorlabs to ask for the data since it's specific to our laser). I have not figured out why there is a shift between the trials- it corresponds to the dips on the signal downfall (groups 2 and 4) but not sure why since the signal is the same at these points. If you zoom in onto the smaller dip, you see two overlapping profiles, potentially from two closely spaced transitions. Will determine what wavelengths these correspond to after getting Thorlabs data. Also, previously when we measured the absorption, we had a laser current of around 203 mA, so it was probably saturating the photodetector.

We mounted and placed the EOM (EO-PM-NR-C1) on the vapor cell setup board. Attached is the scanned sheet it came with and the new schematic since the EOM is now free space (need an additional 780 mirror). We will need lenses for focusing the beam down to the 2mm aperture and possibly a waveplate to make sure the input beam is vertically polarized. Also learned to make BNC cables (https://wiki.mccullerlab.com/DCC/T2400001). 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Daniel Grass - 19:14, Tuesday 16 July 2024 (11750)

Based on the laser specs found here, the spacing between the peaks matches the Rb87 data found here to ~10%.

Lab Infrastructure General
Daniel Grass - posted 18:47, Monday 15 July 2024 (11742)
Oven Baked out at 200°C

I baked out the oven over the weekend at 200°C for 44 hours. It took 104 minutes to reach 199°C from room temperature. I triggered some internal alarm, maybe because I set the maximum temperature too low. This alarm ended the cycle, so I had to start it again. The maximum temperature reached was 204.5°C. The oven tends to oscillate between 197°C and 203°C. 

Surprisingly, the oven vacuum went from 15 Torr to ~1 Torr. I am not sure if this is because the pump was on for a lot longer than previously, or baking out at 200°C allowed the oven to be evacuated more. 1 Torr is much closer to the pressure we should be able to reach based on the pump out speed, leak rate, and minimum pump pressure (0.2 Torr, the minimum pump pressure).

9 hours after turning the heaters off, the oven is 46.7°C. It will probably take another 4 hours to cool. If you want to use an object, I would set the oven to stop heating ~12 hours before you need it.

Lab Infrastructure General
Daniel Grass - posted 18:38, Monday 15 July 2024 (11741)
Ultrasonic Cleaner Setup in the Chemistry Room

I cleaned the fume hood with paper towels and DI water and set up the Vevor Ultrasonic Cleaner (internal designation MX-S24002) in said fume hood. I filled it with 4 L of DI water and turned it on for 30 minutes. I set the heater to 30°C and the temperature reached 41°C (at least). I am not sure why this is, but something to look out for.

GQuEST End Mirror Mount (GQuEST)
Daniel Grass - posted 18:35, Monday 15 July 2024 (11740)
Optics Assembled for Shack–Hartmann Wavefront Sensor

[Erin, Daniel]

Erin and I assembled the optics for the Shack–Hartmann wavefront sensor, except the camera itself. This includes a 780 nm PBS on a Thorlabs KM100PM, a 780 nm quarter waveplate in a RSP1, the mount for end mirror holder, and a beam dump. We will need to figure out a space for this sensor. Right now, the optics are fixed to the table near the assembly area. We also need a few components from Thorlabs and to clean the end mirror mount parts.

GQuEST Output Filter Cavities
Torrey Cullen - posted 18:03, Monday 15 July 2024 (11739)
FC 2 and 3 on one sled

[Erin, Torrey]

We have all optics placed on the table (minus a 775 50:50 BS that needs to be purchased) as a proof of concept of my last post. A few changes have been made since then, mainly not having flip mounts as the input to the cavity. Tomorrow I will start feeding laser light to this side of the table, doing rough alignment, installing any permanent PDs, wiring for flip mounts and PDs, and then finally superoptics.

 

Non-image files attached to this report
Atoms General
Briana Chen - posted 10:32, Sunday 14 July 2024 - last comment - 21:47, Monday 15 July 2024(11736)
Absorption dips :)

[Briana, Ian, Torrey]

We measured absorption dips from sending the 780 nm laser through the vapor cell and measuring the output with a photodetector. The photodetector was connected with a Pomona 2249-C-96 (RG58C/U, 50 Ohm) BNC cable to the smaller Moku Go #01 (previously in the glass cabinet on the right side of the lab upon entrance). We scan through the laser's frequency by changing the laser temperature using the temp controller. Based on the laser diode spec sheet (Figure 3), at a bias current of ~200 mA, the wavelength of 780 nm is between the 25 and 35 degrees Celsius curve (corresponding to 9.99 and 6.53 kOhms respectively assuming the resistance/temperature relationship actually follows the Steinhart equation (it is slightly off but close)), so we manually changed the temperature between 6.53 and 9.99 kOhms until we saw a dip in the power reading. This occurred between 7-8 kOhms so we scanned this range of the temperature controller. We set our bias current to 203.59 mA. 

We first found the appropriate voltage to apply for scanning. From the TED200C temperature control manual, the conversion coefficient for the TH 20K (which we are using) is 2 kOhms/V. For the 1 kOhm range that we want to scan over, this means we need the signal to reach an amplitude of 1 kOhm/(2 kOhms/V) = 0.5 V. So, we manually set the temperature control at 6.996 kOhms (when the output is reset, you restart at this resistance value) and started scanning with a ramp signal (on precision acquisition mode, triggers on the rising edge) at a frequency of 100.000 mHz (determines how long it takes for you to scan the 1 kOhm range?), amplitude of 500.0 mVpp (0.5 V calculation described earlier), offset 250.0 mV (shift up the signal so it begins at 0). This enables scanning from 7-8 kOhms.

We control the set temperature by applying this voltage to the Tune In input (R1) on the back of the TED200C. The input of the temperature control is the tune-in port, which is hooked to the Moku output. The output is hooked up to the Moku input so we can see the signal. You can see consistent dips at a certain temperature, which shows the 780 ish transition is happening! We will use the vapor cell temperature controller to get a deeper dip/better resolution since the temperature of the vapor cell currently fluctuates freely. After plotting data (see attached jupyter notebook), it seems that the ~780 nm transition occurs consistently around 7.5658-7.5690 kOhms. Data is saved under Nextcloud under GQuEST/Vaporcell. Also, I fit these dips to Gaussian/Lorentzian/Voigt profiles, which I think is valid since the time relates to temperature which relates to wavelength so it should have a spectral line shape (dipsnfits for time axis and dipsnfits_temp for conversion to temperature axis). I think the Voigt profile (convolution of Gaussian/Lorentzian) would make the most sense to fit to since there is probably a combination of homogeneous (Lorentzian) and inhomogeneous (Gaussian) broadening in the system. There's not a lot of data points to work with though and the tails have some curve to them because of scanning, so the fits should be taken with a grain of salt. Jupyter notebook has these plots + relevant values. Maybe you can get how much the laser wavelength fluctuates by from the FWHM or standard deviation of temperature at minimum transmittance? (since temperature relates to wavelength) Not sure, should probably think more about that.

Also, we confirmed that the measured power of the 780 nm laser with the black stick-like photodetector agrees with the spec sheet for 25 and 35 celsius at various current settings (160-240 mA) (laserpowerconfirmation.png picture) on 500 mW setting. Relationship between power and current seems linear like the images in the spec sheet. Also, all twelve 780 mirrors for the vapor cell setup have been mounted.

Images attached to this report
Non-image files attached to this report
Comments related to this report
Briana Chen - 21:47, Monday 15 July 2024 (11744)

Update: these data files are now located in the Nextcloud under Users/briana/20240712_Absorption_Meas/

Controls Optimal Controls
Ian MacMillan - posted 16:57, Saturday 13 July 2024 (11738)
Strongly Stable Optimal Controllers

The controllers presented in log post [11731], while leading to a stable closed loop system, are not in themselves stable. The proper term for this is they are not "strongly stable." This does not affect their performance and they are the optimal controller however it may be harder to implement unstable controllers. In this case it may be helpful to find a way to ensure that the final system is strongly stable. There are a few papers on this by various authors on the topic with two being the most useful because they are by authors that we used to write buzz's \(H_2\) /\(H_\infty\) code. The first is \(H_2\) and Mixed  \(H_2\)/ \(H_\infty\) Stable Stabilization (Google Scholar) which lays out the fixed order solution for \(L_2\) disturbances with a fixed \(H_\infty\) performance bound in section 5. The second is Stable Stabilization with \(H_2\) and \(H_\infty\) Performance Constraints (Google Scholar) which gives a more generalized solution for strongly stable controllers. While the second paper says that it is giving a mixed norm result, I'm not sure it is. I would call mixed norm a minimization of \(||H_2|| + ||H_\infty||\) which is diffrent than our goal of minimizing \(||H_2||\) given a \(H_\infty\) bound. Given the papers language and references to its first reference which is about LQG with an \(H_\infty\) bound, I think in reality it is not a mixed norm bit in fact \(||H_2||\) given a \(H_\infty\) bound.

Lab Infrastructure General
Daniel Grass - posted 13:49, Saturday 13 July 2024 (11737)
Vacuum Oven slightly Less Leaky after Tightening Hose around Vacuum Oven Nozzle

[Alex, Torrey, Daniel]

We vented and rotated the vacuum oven so that I could more easily access the back of the oven to tighten the hose clamp around the vacuum oven nozzle. I used (and returned) a ratchet with a Philips attachment from Nick Hutzler's Lab which is very useful and we should buy. We rotated it back, and I evacuated the oven for 30 minutes and it got down to 15 Torr, the same as before tightening. I turned the pump off and kept the valve to the pump open. 20 hours later, the pressure rose to ~28 Torr. This is a slight improvement of the 30 Torr after 20 hours from the previous test. I cannot tighten the hose clamp further because the hose is starting to get shreded. I think 15 Torr is about the best we are going to get with this oven and pump.

I am now going to bake out the oven itself at 200°C for 48 hours. Assuming that goes well, I think the oven is ready to use.

Controls Optimal Controls
Lee McCuller - posted 16:17, Friday 12 July 2024 (11735)
Vega-Altair plotter for optimal controls plots

I've wanted to create an example of how to bundle data and plot it through altair for a while. The recent optical controls work has given a good example to start from. Just pushed SHA d4ebcc2d0a7ed into the dev branch.

This adds a notebook vega_plot_AIC.ipynb that loads the pickle of controllers created by the optimal control scanner. After loading, this object is a list of dictionaries. Each dictionary maps a set of column-keys to values. Some of those values are floats, like the RMS of the figure of merit, or the phase margin, some are strings, like which solver. Some of them are objects (the wield.control.SISO or wield.control.MIMO of the controllers or system) and some of them are arrays of data, like from frequency response functions.

You can then process that list to drop the objects and setup the types of the arrays. You can then load them into a dataframe. I'm using polars as it seem to identify the lists-of-floats well, but pandas might be fine for this. Polars is much more complicated to use, but enforces mannerisms that are very fast.

After loading the data, then you can ship it to a modern plotter like altair. The frequency response functions are very large, so you have to be careful with them, to keep them as arrays. To plot them, you have to "explode", "melt" or "flatten" the arrays into (many) more data columns. Pandas and polars use explode and melt. Altair uses the "transform_flatten". This expands the data so that you can create line plots, but you can't ship expanded data because the resulting json uses too much memory (trust me)

Digression: Some might say I like Altair because it uses the type_action verbiage that my libraries use.

The linked notebook shows the data transform tricks so that this set of plots, around 100MB binary data, can be sent to a browser as JSON and not cause it to grind to a halt. Seems to work, but I haven't been sure.

 

The plots are interactive, you can scroll, zoom, and select points. Getting them to be comparable to Ians matplotlib ones will still be a fair amount of work, but this framework will allow it to happen. Note that altair has very unusual methodologies that are not intuitive, so this is far from a simple example. See their example library at https://altair-viz.github.io/gallery/index.html for more ideas.

We can make selectors based on the H2 vs HiB, or based on iso-zeta lines, and the coloration could use a lot of work (like having the fresponse color on igsq when when iso-zeta is selected). The phase plots could also probably be made to wrap the phase.

Anyway, it's pretty cool. I don't intend to work on this further except to help any of you take it from this demo into a more fully-nice setup.

Hmm - something about the log is causing the html attachement to time-out and fail.

Use the direct download link git download link instead (and then open the local file in browser).

Non-image files attached to this report
Controls Optimal Controls
Ian MacMillan - posted 13:22, Friday 12 July 2024 (11731)
Exporting Optimal Linear Controllers for C-HARD Pitch

I exported the controllers for the C-HARD Pitch DOF from log post [11723]. I attached the exported file (AIC_filters.yml). A controller can be loaded by selecting a desired controller from the RMS plot (AIC_Hib_FOM_RMS_Scatter.pdf) then plugging that index into the python file included (load_K.py) as index_of_interest. This will load the filter as a scipy zpk. The order of these controllers is rather high so they may need to be reduced further for use in a LIGO CDS filter bank or at least broken into multiple filters in series. The code in the load_K.py file has been added to the file AIC_calc_ctrl.ipynb in the buzz/dev repo.

I also attached the requirements.txt file for installing the wield packages. It installs the all the relevant wield packages and not just the needed ones for loading the .yml file.

Non-image files attached to this report
Equipment General
Torrey Cullen - posted 11:42, Friday 12 July 2024 - last comment - 16:02, Friday 12 July 2024(11729)
Location of tools to make homemade shielded BNC
Images attached to this report
Comments related to this report
Ian MacMillan - 12:39, Friday 12 July 2024 (11730)

This holds the record as the shortest logpost.

Lee McCuller - 16:02, Friday 12 July 2024 (11734)

Put quite a keenly useful log!

There is a wire spooler in the new makerspace, that the cable spool should eventually go on. I also have a couple more wire stripping tools to collect.

Atoms General
Briana Chen - posted 21:06, Thursday 11 July 2024 - last comment - 13:00, Friday 12 July 2024(11728)
Preliminary Vapor Cell Setups

[Briana, Ian, Torrey]

We currently have the current controller limit set to 230 mA, but the power sent to the laser doesn't seem to be able to reach the standard 39.4 mW in the specs before the current applied reaches 250 mA. At 236 mA, temperature set at 9.99 kOhms (supposedly 25 deg Celsius), the power reading is 35 mW (with the 500 mW setting in place). This reading agrees with graphs on the spec sheet but when we set the current control limit, we should theoretically be able to turn up the current to some value <250 mA such that the maximum power output (55mW) is reached (see thorlabs video linked two posts back). The lower laser power may be because we aren't driving temperature high enough for the laser to reach the higher power output.

We did use the 5 mW setting on the photodetector at some point and measured 39.4 mW at 230 mA at temperature ~25 deg Celsius. We will not use the 5 mW setting moving forward because our power is much higher than that. The current/corresponding power readings on the previous post should be disregarded becaues we also used the 5 mW setting to measure things, which did not have the ND filter from the 500 mW setting in place. 

We also set up the plan to shoot the 780 nm laser through the vapor cell and scan across the frequency to detect the Lamb dip (beam path in red in one of the pictures). This will be done with a photodiode (PDA36A2) connected to a Moku, which we will set up tomorrow. The temperature control for the vapor (TC300) has also been placed but not set up. We should also check the specs of the vapor cell to see what current is needed.

Images attached to this report
Comments related to this report
Ian MacMillan - 13:00, Friday 12 July 2024 (11733)

I misread the power sensor ND filter and am responsible for that mistake. I did not put enough power/area to get anywhere close to damaging the sensor. We should print a label for the sensor so in future people don't repeat my mistake.

Lab Infrastructure General
Daniel Grass - posted 17:56, Thursday 11 July 2024 (11727)
Vacuum Oven Temperature Change Rate is Working

The Vacuum Oven pretty clearly limited its temperature output when going to 100°C when I limited the change to 1°C/minute. This could be helpful for annealing or baking out substances with low thermal conductivity.

Displaying reports 561-580 of 1091.Go to page Start 25 26 27 28 29 30 31 32 33 End